OpenStudy (nathaliama13):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: C = 5 over 9(F − 32) Which of the following formulas correctly solves for F? f=9/5 c+32 f=9/5 c−32 f=9c + 32/5 f = 9c − 32/5

OpenStudy (fortytherapper):

So is it 5 -------- 9(F - 32) Or 5 ------ (F - 32) 9 ??

OpenStudy (nathaliama13):

Second one

OpenStudy (fortytherapper):

So if we wanna get F by its self, we have to do opposite operations. First things first, how would you distribute the 5/9?

OpenStudy (skullpatrol):

$$\Huge C=\dfrac{5}{9}(F-32)$$ $$\Huge \dfrac{9}{5}C=F-32$$ $$\Huge \dfrac{9}{5}C+32=F$$